Hilda adds 5 to a number, then multiplies the sum by negative 2. The result is 6. Write an equation to find the number, x.
What is the number? Enter your answer in the box.
1
Expert's answer
2020-04-21T11:36:35-0400

We can write an expression in the following form:

"(x+5)\cdot (-2)=6,"

"x+5=-3,"

"x=-8"